I seriously need help with quadratic formulas. please help.


-55x+30=50x2

Found 2 solutions by stanbon, checkley77:
Answer by stanbon(75887)

-55x+30=50x2
---
Rearrange:
50x^2 + 55x - 30 = 0
---
Divide thru by 5 to get:
10x^2 + 11x - 6 = 0
---
a = 10, b = 11, c = -6
----
Formula: x = [-b +- sqrt(b^2-4ac)]/(2a)
---
x = [-11 +- sqrt(121 - 4*10*-6)]/(20)
---
x = [-11 +- sqrt(361)]/(20)
---
x = [-11 + 19]/(20) or x = [-11 - 19]/(20)
---
x = (8/20) or x = -3/2
---
x = 2/5 or x = -3/2
==============================
Cheers,
Stan H.

Answer by checkley77(12844)

-55x+30=50x^2
50x^2+55x-30=0
5(10x^2+11x-6)=0
5(5x-2)(2x+3)=0
5x-2=0
5x=2
x=2/5 ans.
2x+3=0
2x=-3
x=-3/2 ans.